OpenStudy (anonymous):

explain what meant by stating that the laplace transform is a linear transform... write down the two linear properties..

OpenStudy (badhi):

if a transform is a linear transform, $$T(u+v)=T(u)+T(v)\\ T(ru)=rT(u)$$ where r is constants, u,v are vectors so, laplace tansform is a transform of a function defined as $$T[f(t)]=\int_0^\infty f(t)e^{-st}dt$$ in this case u(or v) that is stated previously is a function. Check for the linearity.